# Synthetic Data for Robust Stroke Segmentation

### v2

@@ Abstract
Current deep learning-based approaches to lesion segmentation in neuroimaging often depend on high-resolution images and extensive annotated data, limiting clinical applicability. This paper introduces a novel synthetic data framework tailored for stroke lesion segmentation, expanding the SynthSeg methodology to incorporate lesion-specific augmentations that simulate diverse pathological features. Using a modified nnUNet architecture, our approach trains models with label maps from healthy and stroke datasets, facilitating segmentation across both normal and pathological tissue without reliance on specific sequence-based training. Evaluation across in-domain and out-of-domain (OOD) datasets reveals that our method matches state-of-the-art performance within the training domain and significantly outperforms existing methods on OOD data. By minimizing dependence on large annotated datasets and allowing for cross-sequence applicability, our framework holds potential to improve clinical neuroimaging workflows, particularly in stroke pathology. PyTorch training code and weights are publicly available at https://github.com/liamchalcroft/SynthStroke, along with an SPM toolbox featuring a plug-and-play model at https://github.com/liamchalcroft/SynthStrokeSPM.
@@ 1 Introduction
Dense classification of images or volumes, known as semantic segmentation, is crucial in many neuroimaging and medical imaging analysis pipelines. Properties such as volume and surface distance of healthy tissue structures are used in longitudinal models to measure anatomical changes, while tumor labels are essential for radiotherapy planning. Extracting fine-grained labels has typically been limited to high-contrast, high-resolution structural scans, such as those acquired using the MPRAGE sequence, which are rarely found in clinical settings. This limitation poses a significant barrier to applying these methods in clinical practice.
@@
Prior work to SynthSeg demonstrated the potential of encoding anatomical priors in a neural network through pre-training with unpaired parcellation labels (Dalca et al., 2018). Such methods face significantly larger challenges when applied to the heterogeneous shape and spatial distribution of lesions. In healthy parcellation, anatomical structures have consistent positions across individuals (e.g., the brainstem reliably appears in the same region of the brain), a regularity that has motivated atlas-based approaches to parcellation. In contrast, lesions are highly variable across individuals—not only in number and size but also in their spatial distribution. Unlike anatomical structures, lesions cannot be reliably mapped to a specific location or shape within an atlas. Although the exact site of lesion initiation is often influenced by the brain’s vascular architecture—meaning that certain regions are statistically more susceptible to stroke due to the location of large blood vessels—the resulting lesion’s size, shape, and spread are highly variable. Multiple sclerosis serves as an exception, with lesions that are somewhat predictable in their white matter localisation (Lassmann, 2018), enabling modelling through synthetic deep learning frameworks (Billot et al., 2021) and traditional probabilistic models (Cerri et al., 2021). More recently, Liu et al. (2024) has shown promising results by training a SynthSeg-like model on lesion labels from various pathologies, providing a foundation for fine-tuning on multiple downstream datasets. However, achieving a model that can seamlessly generalise for open-domain stroke segmentation remains an active area of research
In our work, we extend the SynthSeg framework to the task of stroke lesion segmentation via a novel lesion-pasting method that better simulates variety in lesion appearance. We show that this can be used to expand an existing segmentation dataset to attain competitive performance in-domain, and state-of-the-art performance out-of-domain that can be further enhanced with existing test-time domain adaptation techniques. We validate this on a comprehensive range of lesion datasets with a wide distribution of image characteristics and lesion physiology. To assist in widespread evaluation of this framwork, we release PyTorch training code/weights, and a MATLAB toolbox for SPM to reduce the barrier to clinical adoption.
@@ 2 Methods
The core premise of our training procedure is to generate synthetic images with paired labels by combining random permutations of healthy tissue labels and binary stroke lesion maps.
A number of important modifications were made from the original SynthSeg procedure to better suit it to the task of stroke segmentation. First, as we are ultimately only interested in annotating stroke infarct and not the surrounding anatomy, we can opt for a much simpler set of labels than the full set of anatomical labels generated from Freesurfer in SynthSeg. For our experiments we opt for the posterior maps generated using the MultiBrain toolbox in SPM (Brudfors et al., 2020). This reduces the number of sampled healthy labels to only 9 classes, and requires using a single model to generate labels instead of the two used in SynthSeg. A comparison of the sampled labels from Freesurfer and MultiBrain can be seen in Figure 1. Sampled labels are summed using a weighting based on their posterior likelihoods, allowing for the simulation of partial volume. Of these classes, only 4 are labels of brain tissue (Gray Matter (GM), White Matter (WM), GM/WM partial volume and Cerebrospinal Fluid (CSF)) which also makes for a lower memory footprint in training. In all Synth-model training, we opt for predicting these 4 tissue classes in addition to the lesion class in order to prevent issues of class imbalance that can occur when treating lesion segmentation as a binary task.
@@
We also enhance the default training procedure by introducing lesion-specific augmentations to mimic variability in lesion appearance. We first use the Soft-CP algorithm (Dai et al., 2022) with random dilations and erosions to vary how the infarct is blended into its neighbouring healthy tissue. We also introduce a spatially varying multiplicative field to the binary lesion map, which mimics the variety in the local texture of a lesion and presence of penumbra within a stroke. This is performed by applying the MONAI Random Bias Field augmentation (Cardoso et al., 2022) and thresholding the result in the $[0,1]$ range. A similar approach was adopted in Middleton et al. (2024) and showed in improvement in model sensitivity.
A model was trained using a combination of healthy labels from the OASIS-3 dataset (LaMontagne et al., 2019) and lesion masks from the ATLAS dataset (Liew et al., 2022). In the OASIS-3 dataset (N=2679, 2579:100 train-val split), tissue labels were extracted using SPM’s MultiBrain (Brudfors et al., 2020) and spatially aligned to ICBM space to enable correct lesion placement. The ATLAS dataset (N=655, 419:105:131 train-val-test split) was used for pasting lesion maps, which were also aligned to ICBM space. A lesioned brain map was generated by randomly pasting an ATLAS lesion label onto an OASIS subject’s parcellation using the previously described augmentations.
The generated maps were used to randomly assign Gaussians to each tissue class by sampling a mean from the distribution $\mu\sim\text{Uniform}(0,255)$, a standard deviation $\sigma\sim\text{Uniform}(0,16)$ and the FWHM of the applied smoothing from $\text{Uniform}(0,2)$. All images were then extensively augmented using a variety of transformations. These include the application of a random bias field, which involves rescaling intensities based on a smooth multiplicative field controlled by a number of control points $\mathcal{U}(2,7)$ and a bias strength $\mathcal{U}(0,0.5)$, with multiplicative values sampled from $\mathcal{U}(1-\text{strength},1+\text{strength})$. Random affine deformation is also employed, consisting of random rotations within $\mathcal{U}(-15,15)$ degrees, shearing within $\mathcal{U}(0,0.012)$, and zooming within $\mathcal{U}(0.85,1.15)$. Additionally, random elastic deformations are applied with a maximum displacement $\mathcal{U}(0,0.05)$ using control points $\mathcal{U}(0,10)$. Another transformation simulates random skull-stripping by generating a perfect foreground mask of brain tissues and introducing flaws through random dilation (probability of 0.3, radius of 2) and erosion (probability of 0.3, radius of 4). Random flips are applied across each axis with a probability of 0.8. Random Gaussian noise is introduced, controlled by a signal-to-noise ratio (SNR) $\mathcal{U}(0,10)$, and smoothed by a g-factor field with smoothness $\mathcal{U}(2,5)$. Slice anisotropy is simulated by rescaling the slice thickness based on a resolution factor $\mathcal{U}(1,8)$, with experiments assuming an original isotropic resolution of 1mm3. Random contrast adjustment is applied using a gamma factor $\gamma$ sampled from $10^{\mathcal{N}(0,0.6)}$, while random motion blurring is simulated using a point-spread function with a full width at half maximum (FWHM) $\mathcal{U}(0,3)$. Intensities were clipped to the 1st and 99th percentiles and images z-normalised to a zero mean and unit standard deviation. Training was performed at an isotropic spacing of 1 mm and volumes randomly cropped to 192$\times$192$\times$192. A simplified schematic (with image-quality augmentations removed) is shown in Figure 2.

### v3

@@ Abstract
Current deep learning-based approaches to lesion segmentation in neuroimaging often depend on high-resolution images and extensive annotated data, limiting clinical applicability. This paper introduces a novel synthetic data framework tailored for stroke lesion segmentation, expanding the SynthSeg methodology to incorporate lesion-specific augmentations that simulate diverse pathological features. Using a modified nnUNet architecture, our approach trains models with label maps from healthy and stroke datasets, facilitating segmentation across both normal and pathological tissue without reliance on specific sequence-based training. Our method achieves robust out-of-domain performance where conventional approaches fail, with in-domain performance of 48.2% Dice compared to 57.5% for conventional training. Crucially, even with oracle knowledge of the optimal domain adaptation method - an unrealistic scenario in practice - conventionally-trained models cannot match our synthetic approach in out-of-domain settings. The framework demonstrates that synthetic pre-training provides fundamental robustness unachievable through test-time adaptation alone. Our approach reduces reliance on domain-specific training data and helps bridge the gap between research-grade and clinical scans to improve clinical stroke neuroimaging workflows. PyTorch training code and weights are publicly available at https://github.com/liamchalcroft/SynthStroke, along with an SPM toolbox featuring a plug-and-play model at https://github.com/liamchalcroft/SynthStrokeSPM.
doi:
https://doi.org/10.59275/j.melba.2025-f3g6
@@ 1 Introduction
Semantic segmentation is a critical component of neuroimaging pipelines, enabling precise quantification of anatomical structures and lesions for applications like tracking disease progression and planning treatments. In research settings, segmentation labels are typically derived from high-quality, standardised structural scans (e.g., MPRAGE) that benefit from consistent field-of-view, spacing, orientation, and minimal artifacts. In contrast, clinical scans often exhibit significant variability in these factors, which can severely impact deep learning model performance. Consequently, models trained on homogeneous research-grade data may not generalise well to the diverse, lower-quality images encountered in clinical practice.
@@
Prior work to SynthSeg demonstrated the potential of encoding anatomical priors in a neural network through pre-training with unpaired parcellation labels (Dalca et al., 2018). Such methods face significantly larger challenges when applied to the heterogeneous shape and spatial distribution of lesions. In healthy parcellation, anatomical structures have consistent positions across individuals (e.g., the brainstem reliably appears in the same region of the brain), a regularity that has motivated atlas-based approaches to parcellation. In contrast, lesions are highly variable across individuals - not only in number and size but also in their spatial distribution. Unlike anatomical structures, lesions cannot be reliably mapped to a specific location or shape within an atlas. Although the exact site of lesion initiation is often influenced by the brain’s vascular architecture - meaning that certain regions are statistically more susceptible to stroke due to the location of large blood vessels - the resulting lesion’s size, shape, and spread are highly variable. Multiple sclerosis serves as an exception, with lesions that are somewhat predictable in their white matter localisation (Lassmann, 2018), enabling modelling through synthetic deep learning frameworks (Billot et al., 2021; Laso et al., 2024) and traditional probabilistic models (Cerri et al., 2021). More recently, Liu et al. (2024) has shown promising results by training a SynthSeg-like model on lesion labels from various pathologies, providing a foundation for fine-tuning on multiple downstream datasets.
Robust open-domain stroke segmentation remains an unsolved challenge. Most domain-specific frameworks for stroke lesion segmentation are targeted towards lesion-pasting (Zhang et al., 2021; Dai et al., 2022; Basaran et al., 2023), aiming to augment the anatomical variety without making any attempt to augment the variety in image contrasts. Likewise, label-conditioned generative models such as TumorGAN (Li et al., 2020) can similarly only generate new lesioned brains within the learned distribution of image intensities. None of these works approach the task of robustness to shifts in image appearance, instead focusing on shape-related augmentation.
In our work, we extend the SynthSeg framework to the task of stroke lesion segmentation via a novel lesion-pasting method that better simulates variety in lesion appearance. Our hybrid approach trades a statistically significant 9.3% median Dice reduction in-domain (57.5% vs 48.2%, p<<0.001) for improved out-of-domain robustness. Crucially, we demonstrate that even with oracle knowledge of the optimal domain adaptation method, conventional training cannot match our synthetic approach in out-of-domain scenarios. We validate this on a comprehensive range of lesion datasets with a wide distribution of image characteristics and lesion physiology. To assist in widespread evaluation of this framework, we release PyTorch training code/weights, and a MATLAB toolbox for SPM to reduce the barrier to clinical adoption.
@@ 2 Methods
Terminology and Notation
For clarity, we define the key terms used throughout this work:
  • TTA (Test-Time Augmentation): A procedure at inference where multiple augmented versions (here generated by flips) of an input are processed and their predictions averaged to improve robustness.
  • DA (Domain Adaptation): Techniques applied at test time to adapt a trained model to new data distributions.
  • Oracle DA: The hypothetical best-case scenario where the optimal DA method is known a priori for each dataset/modality combination.
2.1 Synthetic Data Generation Framework
Rationale.
Our goal is to create a large, diverse and perfectly labelled training set without the labour of voxel-wise annotation. We therefore generate paired image-label volumes by composing healthy tissue maps with realistically shaped stroke lesions, followed by intensity synthesis and heavy image-quality augmentation (Fig. 2).
(i) Healthy-tissue label bank.
Instead of the 100+ FreeSurfer classes used by SynthSeg, we adopt the nine posterior tissue maps produced by MultiBrain (Brudfors et al., 2020). This reduces memory usage, speeds up sampling and still retains the GM/WM/CSF boundaries that matter for lesion realism (Fig. 1).
@@ 2.1 Synthetic Data Generation Framework
(ii) Lesion Copy-Paste (Lesion-CP).
We extend Soft-CP (Dai et al., 2022) with random dilate/erode ’feathering’ and a spatially varying bias-field multiplier (MONAI Random Bias Field (Cardoso et al., 2022)) to mimic penumbral intensity fall-off (Middleton et al., 2024).
(iii) Intensity sampling.
Each tissue class is assigned $\mu\!\sim\!\mathcal{U}(0,255)$, $\sigma\!\sim\!\mathcal{U}(0,16)$ and Gaussian blur $\text{FWHM}\!\sim\!\mathcal{U}(0,2)$. For stroke lesions we modulate the copy-pasted mask with the bias field to create intra-lesion heterogeneity. We sample from a single Gaussian distribution, with the implications of this choice examined via Hartigan’s dip-test in Appendix A.
(iv) Image-quality augmentations.
Table 1 summarises every random transform (bias field, affine, elastic, skull-strip imperfection, noise, resolution, motion, contrast, etc.).
Future work will explore mixture-of-Gaussians modelling (Ashburner and Friston, 2005) to represent lesions that are simultaneously hyper- and hypo-intense.
Table 1: Parameter ranges for every image-quality augmentation used during synthetic training. indicates augmentations applied only to synthetic data and not to real ATLAS samples.
Category Transform Sampling range / notes
Bias field Multiplicative bias Control points $\mathcal{U}(2,7)$, strength $\mathcal{U}(0,0.5)$
Affine Rotation $\mathcal{U}(-15^{\circ},15^{\circ})$ on each axis
Shear $\mathcal{U}(0,0.012)$
Zoom $\mathcal{U}(0.85,1.15)$
Elastic Deformation grid Control points $\mathcal{U}(0,10)$, max disp. $\mathcal{U}(0,0.05)$
Skull-strip flaw Dilation $p=0.3$, radius = 2 vox.
Erosion $p=0.3$, radius = 4 vox.
Noise Gaussian (SNR) SNR $\mathcal{U}(0,10)$, smoothed by $g$-factor $\mathcal{U}(2,5)$
Resolution Slice anisotropy Thk. factor $\mathcal{U}(1,8)$ (base res. 1 mm3)
Contrast Gamma $\gamma=10^{\mathcal{N}(0,0.6)}$
Motion blur PSF width (FWHM) $\mathcal{U}(0,3)$ vox.
Flip Mirror on each axis $p=0.8$